# Retrieval-Augmented Generation for AI-Generated Content: A Survey

### v2

@@ III-A2 Latent Representation-based RAG
In the framework of Latent Representation-based RAG, the generative models interact with latent representations of retrieved objects, thereby enhancing the model’s comprehension abilities and the quality of the content generated.
@@
The FiD [34] technique leverages both BM25 and DPR for sourcing supportive paragraphs. It concatenates each retrieved paragraph and its title with the query, processing them individually through the encoder. FiD reduces computational complexity and efficiently utilizes relevant information to generate answers by fusing information from multiple retrieved paragraphs in the decoder, rather than processing each paragraph in the encoder. The application of Fusion-in-Decoder methodologies transcends the realm of textual content processing, demonstrating substantial potential and adaptability in processing code, structured knowledge, and diverse multimodal datasets. Specifically within the code-related domain, technologies such as EDITSUM [139], BASHEXPLAINER [140], and RetrieveNEdit [141] adopt the FiD approach, facilitating integration through encoder-processed fusion. Re2Com [142], and RACE [143] , among other methods, also feature the design of multiple encoders for different types of inputs. In the field of Knowledge Base Question Answering (KBQA), the FiD method has been widely adopted, demonstrating significant effectiveness. UniK-QA [144], DECAF [145], SKP [146], KD-CoT [147], and ReSKGC [148] have effectively enhanced the performance of QA systems through the application of Fusion-in-Decoder technology. This illustrates that by integrating RAG for KBQA, the efficiency and accuracy of QA systems can be significantly improved. In the field of Science, RetMol [54] employ the Fusion-in-Decoder strategy, integrating information at the decoder stage to enhance the relevance and quality of the generated molecular structures.
@@
Retro [35] pioneers the integration of retrieved text via ”Chunked Cross-Attention,” a novel mechanism that segments the input sequence into discrete chunks. Each chunk independently executes cross-attention operations, thereby mitigating computational burdens. This technique enables the model to selectively retrieve and assimilate distinct documents for varied sequence segments, fostering dynamic retrieval throughout the generation process. This enhances the model’s adaptability and enriches the contextual backdrop of generated content. In the domain of image generation, cross-attention mechanisms have been widely adopted within RAG frameworks. Methods such as Re-imagen [149], KNN-Diffusion [150], RDM [151] and LAION-RDM & ImageNet-RDM [152] utilize cross-attention to integrate multiple retrieval results, effectively enhancing the overall performance of the models. On the other hand, Li [153] introduces the ACM, a text-image affine combination module, which notably does not employ any form of attention mechanism.
@@
Diverging from prior methods for knowledge, EaE [154] empowers language models to internalize explicit entity knowledge. EaE introduces an entity-specific parameterization, optimizing inference efficacy through an entity memory layer embedded within the transformer architecture. This layer directly acquires entity representations from textual data, utilizing a sparse retrieval strategy to fetch the nearest entities based on their embeddings, thus refining the model’s comprehension through a calculated aggregation of entity-specific information. Distinctly, TOME [155] shifts the focus towards comprehensive mention encodings, prioritizing the granularity of mention over mere entity representations. It meticulously generates encodings for each entity mention across Wikipedia, populating a repository with approximately 150 million entries. This repository, encompassing key and value encodings alongside entity IDs, empowers the retrieval of much more fine-grained information. TOME integrates an initial transformer block to process input texts, followed by TOME blocks with memory attention layers, facilitating the synthesis of multifaceted information sources and enhancing inferential reasoning capabilities, even for unencountered entities. Memorizing Transformers [30] revolutionize long document processing through the integration of a kNN-augmented attention mechanism within a Transformer layer. This innovation triggers a kNN search amidst input sequence processing, fetching data based on similarities between the sequence and stored key-value pairs, thereby elevating performance without necessitating complete retraining. This approach not only bolsters processing efficiency but also broadens the model’s memory span, enabling self-retrieval from its generated outputs and fine-tuning for extensive knowledge bases or code repositories. Unlimiformer [156], by embedding a k-nearest neighbors (kNN) index within a pre-trained encoder-decoder transformer framework, pioneers handling inputs of indefinite length. Storing hidden states of input tokens in the kNN index allows for the efficient retrieval of highly relevant tokens during decoding. This innovation extends the model’s capacity to manage prolonged sequences.
@@
In the field of 3D generation, ReMoDiffuse [50] introduces a semantics-modulated attention mechanism. This technology enhances the accuracy of generating corresponding 3D motions based on textual descriptions. AMD [157] achieves efficient conversion from text to 3D motion by fusing the original diffusion process with the reference diffusion process.
@@
In the audio domain, Koizumi [42] utilizes a pre-trained large-scale language model, incorporating dense features generated by VGGish and embedding networks in the attention module to guide the generation of audio captions. Re-AudioLDM [158] extracts deep features from text and audio using T5 and AudioMAE, and integrates these features within the attention mechanism of its Latent Diffusion Model (LDM).
@@
In the field of video captioning, R-ConvED [47] employs a convolutional encoder-decoder network architecture, which processes retrieved video-sentence pairs with the aid of an attention mechanism to generate hidden states and subsequently produce captions. CARE [159] introduces a concept detector as an input to the decoder and embeds concept representations within a hybrid attention mechanism. EgoInstructor [48] employs gated-cross attention to integrate these textual inputs with encoded video features, enhancing the relevance and coherence of the generated captions for egocentric video content.
@@ III-A3 Logit-based RAG
In Logit-based RAG, generative models combine retrieval information through logits during the decoding process. Typically, the logits are summed or combined through models to produce the probability for step-wise generation.
@@
The kNN-LM [36] model integrates a pre-trained neural language model with the k-nearest neighbor search. It employs the pre-trained model to generate a list of candidate words and their probability distribution, while simultaneously performing retrieval from a data repository to find the k most relevant neighbors based on the current context, thus enhancing the output of the original language model. The innovation at the core of this model lies in its ability for dynamic retrieval of information from a broad text corpus, significantly improving the accuracy and relevance of its predictions, particularly in addressing rare patterns and adapting to various fields. He [37] introduces a new framework that is predicated on performing retrieval operations only when necessary, aimed at enhancing the inference efficiency of the kNN-LM model through an adaptive retrieval. This framework accelerates the model’s inference speed by training a retrieval adapter, which automatically identifies and eliminates unnecessary retrieval actions in certain scenarios. This method allows the model to dynamically decide on the necessity of retrieval based on the current context, thereby balancing the trade-off between performance and efficiency, and substantially increasing inference speed while maintaining model performance.
@@
Unlike previous methods that only merge memories during the testing time, TRIME [160] achieves memory merging during both training and testing phases, treating in-batch examples as accessible memory. TRIME leverages new data batching and memory construction techniques to effectively utilize external memory. TRIME employs BM25 scores to pack paragraphs with high lexical overlap into the same batch, constructing the training memory to further optimize model performance. NPM [161] is a non-parametric masked language model comprised of an encoder and a reference corpus. Unlike traditional models that apply a softmax over a finite vocabulary, NPM models a non-parametric distribution over the corpus. The encoder’s role is to map phrases from the corpus into fixed-size vectors, filling in [MASK] by retrieving the phrase most similar to the masked position.
Beyond the text domain, other modalities such as code and image also utilize methods that integrate retrieved content with language models in the final output stage.
@@
For code-to-text conversion task, Rencos [120] generates multiple summary candidates in parallel from the retrieved code. It then normalizes these candidates using edit distance and calculates the final probability to select the summary output that best matches the original code. In code summarization task, EDITSUM [139] enhances the quality of summary generation by integrating prototype summaries at the probability level. For text-to-code tasks, the kNN-TRANX [162] model employs a combination of a confidence network and meta-knowledge to merge retrieved code fragments. It utilizes a seq2tree structure to generate target code that closely matches the input query, thereby increasing the accuracy and relevance of code generation.
In image captioning tasks, MA [163] combines an attention-based encoder-decoder, using the image encoder to extract visual features to construct the semantic part, and decodes it word by word with the information retrieved. MA interpolates between two distributions generated by the caption decoder and the memory-augmented module to determine the distribution of the next word.
@@ III-A4 Speculative RAG
Speculative RAG looks for opportunities to use retrieval instead of generation to save resources and accelerate response speed. REST [31] replaces the small models in speculative decoding [164] with retrieval, so as to generate drafts. GPTCache [38] tries to solve the problem of high latency when using the API of LLMs by building a semantic cache for storing LLM responses. COG [165] decomposes the process of text generation into a series of copy-and-paste operations, which retrieves words or phrases from the retrieved documents instaed of generation. Cao et al. [166] proposes a new paradigm to eliminate the dependence of the final result on the quantity and quality of the first stage retrieve content, which replaces generation by directly retrieving phrase level content.
@@ III-B RAG Enhancements
In this section, we will introduce methods to enhance RAG performance. We categorize existing methods into 5 distinct groups based on their enhancement targets: input, retriever, generator, result, and the entire pipeline.
@@ III-B1 Input Enhancement
The input refers to the user’s query, which is initially fed into the retriever. The quality of the input significantly impacts the final outcome of the retrieval stage. Consequently, enhancing the input becomes essential. In this section, we introduce two methods: query transformation and data augmentation.
@@
Query transformation can enhance the result of retrieval by modifying the input query. Query2doc [167] and HyDE [168] first use query to generate a pseudo document, and then use this document as a key for retrieval. The advantage of doing so is that the pseudo document will contain more rich relevant information, which helps to retrieve more accurate results. TOC [169] employs recursive retrieval augmented clarification on ambiguous questions to construct a tree of disambiguated questions, thereby generating comprehensive answers to these ambiguous questions. During the construction process of this tree structure, TOC utilizes a self-verification pruning method to ensure the factual relevance of each node.
@@
Data augmentation refers to improving the data in advance before retrieval, such as removing irrelevant information, eliminating ambiguity, updating outdated documents, synthesize new data, etc. which can effectively improve the performance of the final RAG system. MakeAnAudio [43] uses captioning and audio-text retrieval to generate caption for language-free audio to mitigate data sparsity, and adds random concept audio to improve the original audio. LESS [170] strategically selects an optimal dataset for downstream tasks by analyzing gradient information, with the aim of maximizing the dataset’s impact on fine-tuning the model’s performance in response to instructional prompts.
@@ III-B2 Retriever Enhancement
In the RAG system, the process of retrieval is crucial. Generally, the better the content quality, the easier it is to stimulate the ability of LLMs in-context learning as well as other generators and paradigms. The worse the content quality, the more likely it is to cause model hallucinations. Therefore, in this section, we will discuss how to efficiently improve the effectiveness of the retrieval process.
Recursive Retrieve
Recursive retrieve is the process of splitting a query before retrieval, and performing multiple searches to retrieve more and higher quality content. Jagerman et al. [171], RAT [172] and ReACT [173] all use Chain-of-Thought(COT) [174] to enable the model to break queries down gradually and provide more rich related knowledge. LLMCS [175] applies this technology to the conversational system and obtained better retrieval results by rewriting the conversation record. ACTIVERAG [176] employs the Cognition Nexus mechanism to skillfully merge retrieved knowledge with the language model’s inherent cognitive system, thus optimizing the COT process and effectively aiding in answer generation. RATP [177] leverages Monte-Carlo Tree Search technique for multi-stage decision-making, where each node represents a thought. The system directs action choices based on scoring model feedback on these thoughts. In node expansion, it merges retrieved documents and thoughts via prompts into the language model to generate new thoughts. Kuratov et al. [178] integrates the Transformer with an RNN, utilizing the model’s intermediate outcomes as the content for retrieval. This process was executed at each layer of the Transformer, thereby significantly extending the text window’s length and effectively mitigating the issue of ”Lost in the Middle”[179].
@@
Chunk optimization technique refers to adjusting the size of the chunk to achieve better retrieval results. Sentence-window retrieval [180] is an efficient approach that enhances retrieval by fetching small chunks of text and returning a window of relevant sentences surrounding the retrieved segment. This method ensures that the context before and after the targeted sentence is included, providing a more comprehensive understanding of the retrieved information. Auto-merge retrieval is another advanced RAG method of LlamaIndex [180] which organizes the document in a tree-like structure, with the parent node containing the content of all children nodes. For example, articles and paragraphs, as well as paragraphs and sentences, all follow a parent-child relationship. In the retrieve process, fine-grained search for children nodes ultimately returns the parent node, effectively providing richer information. To address the lack of contextual information, RAPTOR [181] employs recursive embedding, clustering, and summarization of text chunks until further clustering becomes infeasible, thereby constructing a multi-level tree structure.
Finetune Retriever
@@
As a core component in the RAG system, the retriever plays a crucial role in the entire system operation process. A good embedding model can bring semantically similar content closer together in vector space. The stronger the retriever’s ability, the more useful information can be provided for the subsequent generator, thereby improving the effectiveness of the RAG system. Therefore, the ability of the embedding model [182, 183, 184, 185] is crucial for the overall effectiveness of the RAG system.
@@
In addition, for embedding models that already have good expression power, we can still finetune them using high-quality domain data or task related data to improve their performance in specific domains or tasks. REPLUG [128] treats LM as a black box and update the retriever model based on the final results. APICoder [130] finetunes the retriever with python files and api names, signature, description. EDITSUM [139] finetunes the retriever to decrease the jaccard distance between summaries after retrieval. SYNCHROMESH [121] adds tree distance os ASTs in the loss and uses Target Similarity Tuning to finetune the Retriever. R-ConvED [47] finetunes the Retriever with the same data as generator. Kulkarni et al. [186] applies infoNCE loss to finetune the Retriever.
Hybrid Retrieve
@@
Hybrid retrieve denotes the concurrent employment of a diverse array of retrieval methodologies or the extraction of information from multiple distinct sources. RAP-Gen [187] and ReACC [133] use both dense retriever and sparse retriever to improve the quality of retrieval. Rencos [120] uses sparse retriever to retrieve similar code snippets on syntactic-level and usse dense retriever to retrieve similar code snippets on semantic-level. BASHEXPLAINER [140] first uses dense retriever to capture semantic information and then uses sparse retriever to acquire lexical information. RetDream [49] first retrieves with text and then retrieves with the image embedding. CRAG [188] has designed a retrieval evaluator to assess the relevance of retrieved documents to the input query, triggering three types of retrieval actions based on varying confidence levels: if deemed correct, it directly uses the retrieval results for Knowledge Refinement; if incorrect, it resorts to Web Search; and if ambiguous, it combines both approaches. To enhance performance in question-and-answer tasks, Huang et al. [189] introduces two metrics, DKS(Dense Knowledge Similarity) and RAC(Retriever as Answer Classifier), during the retrieval phase. These metrics account for both the pertinence of the answers and the applicability of the underlying knowledge. UniMS-RAG [190] introduces a novel kind of token, termed as the ’acting token’, which determines the source from which to retrieve information.
@@
The Rerank technique refers to reordering the retrieved content in order to achieve greater diversity and better results. Re2G [191] applies a re-ranker [192] model after the traditional retriever. The effect of the re-ranker model is to re-rank retrieved documents, the purpose of which is to reduce the impact of information loss caused by compressing text into vectors on the quality of retrieval. AceCoder [193] rerank the retrieved programs with a selector. The purpose of introducing the selector is to reduce redundant programs and obtain diverse retrieved programs. XRICL [194] uses a distillation-based exemplar reranker after retrieval. Rangan [195] employs the Quantized Influence Measure, assessing statistical biases between a query and a reference to evaluate the similarity of data subsets, and uses this assessment to rerank retrieval results. UDAPDR [196] utilizes large language models for the mass generation of synthetic queries at low cost, using these synthetic query sets to train domain-specific rerankers. These rerankers act as multiple teachers, and through multi-teacher knowledge distillation techniques, they create a single retriever. LLM-R [197] iteratively trains the retriever, using feedback from a frozen LLM to rank candidate documents and train the reward model. The retriever is further trained based on knowledge distillation. Each iteration of training builds upon the retriever trained in the previous cycle, facilitating iterative optimization in subsequent rounds.
@@
FILCO [198] effectively filters out irrelevant content from the retrieved text chunk, leaving only the precise supporting content. This process simplifies the task for the generator, making it easier to predict the correct answer. FiD-Light [199] initially employs an encoder to convert the retrieved content into a vector, which it then compresses, resulting in a substantial reduction of latency time. RRR [200] integrates the current query with the top-k document in each round through a template, and subsequently restructures it via a pre-trained LLMs(GPT-3.5-Turbo et al.).
@@
In addition to the above optimization methods, there are also some other optimization methods for the retrieve process. For example, Meta-data filtering  [201] is a method to help processing retrieved documents which uses metadata (such as time, purpose, etc.) to filter the retrieved documents for better results. GENREAD [202] and GRG [203] introduces a novel approach where the retrieval process is supplanted or improved by prompting a LLM to generate documents in response to a given question.

### v3

@@ III-A2 Latent Representation-based RAG
In Latent Representation-based RAG framework, retrieved objects are incorporated into generative models as latent representations. This enhances the model’s comprehension abilities and improves the quality of the generated content.
@@
The Fusion-in-Decoder (FiD) [35] technique leverages both BM25 and DPR for sourcing supportive paragraphs. It concatenates each retrieved paragraph and its title with the query, processing them individually through the encoder. FiD reduces computational complexity and efficiently utilizes relevant information to generate answers by fusing information from multiple retrieved paragraphs in the decoder, rather than processing each paragraph in the encoder. The application of Fusion-in-Decoder methodologies transcends the realm of textual content processing, demonstrating substantial potential and adaptability in processing code, structured knowledge, and diverse multimodal datasets. Specifically within the code-related domain, technologies such as EDITSUM [138], BASHEXPLAINER [139], and RetrieveNEdit [140] adopt the FiD approach, facilitating integration through encoder-processed fusion. Re2Com [141], and RACE [142] , among other methods, also feature the design of multiple encoders for different types of inputs.
In the field of Science, RetMol [55] also employs the Fusion-in-Decoder strategy, integrating information at the decoder stage to enhance the relevance and quality of the generated molecular structures.
In the field of Knowledge Base Question Answering (KBQA), the FiD method has been widely adopted, demonstrating significant effectiveness. UniK-QA [143], DECAF [144], SKP [145], KD-CoT [146], and ReSKGC [147] have effectively enhanced the performance of QA systems through the application of Fusion-in-Decoder technology. This illustrates that by integrating RAG for KBQA, the
@@
Retro [36] pioneers the integration of retrieved text via “Chunked Cross-Attention” a novel mechanism that segments the input sequence into discrete chunks. Each chunk independently executes cross-attention operations, thereby mitigating computational burdens. This technique enables the model to selectively retrieve and assimilate distinct documents for varied sequence segments, fostering dynamic retrieval throughout the generation process. This enhances the model’s adaptability and enriches the contextual backdrop of generated content. In the domain of image generation, cross-attention mechanisms have been widely adopted within RAG frameworks. Methods such as Re-imagen [148], KNN-Diffusion [149], RDM [150] and LAION-RDM & ImageNet-RDM [151] utilize cross-attention to integrate multiple retrieval results, effectively enhancing the overall performance of the models.
@@
In addition, there are also some other novel structures worth our attention, Li [152] introduced the ACM, a text-image affine combination module, which notably does not employ any form of attention mechanism. Memorizing Transformers [31] revolutionize long document processing through the integration of a kNN-augmented attention mechanism within a Transformer layer. This innovation triggers a kNN search amidst input sequence processing, fetching data based on similarities between the sequence and stored key-value pairs, thereby elevating performance without necessitating complete retraining. This approach not only bolsters processing efficiency but also broadens the model’s memory span, enabling self-retrieval from its generated outputs and fine-tuning for extensive knowledge bases or code repositories. Unlimiformer [153], by embedding a k-nearest neighbors (kNN) index within a pre-trained encoder-decoder transformer framework, pioneers handling inputs of indefinite length. Storing hidden states of input tokens in the kNN index allows for the efficient retrieval of highly relevant tokens during decoding. This innovation extends the model’s capacity to manage prolonged sequences. Kuratov et al. [154] integrated Transformer with RNN, utilizing the model’s intermediate output as the content for retrieval. This process was executed at each layer of the Transformer, thereby significantly extending the text window’s length and effectively mitigating the issue of “Lost in the Middle”[155]. Diverging from prior methods for knowledge, EaE [156] empowers language models to internalize explicit entity knowledge. EaE introduces an entity-specific parameterization, optimizing inference efficacy through an entity memory layer embedded within the transformer architecture. This layer directly acquires entity representations from textual data, utilizing a sparse retrieval strategy to fetch the nearest entities based on their embeddings, thus refining the model’s comprehension through a calculated aggregation of entity-specific information. on this basis , TOME [157] shifts the focus towards comprehensive mention encodings, prioritizing the granularity of mention over mere entity representations. It meticulously creates a database that stores key and value encodings along with entity IDs, enabling the retrieval of much more fine-grained information. TOME integrates an initial transformer block to process input texts, followed by TOME blocks with memory attention layers, facilitating the synthesis of multifaceted information sources and enhancing inferential reasoning capabilities, even for unencountered entities.
@@
In the field of 3D generation, ReMoDiffuse [51] introduces a semantics-modulated attention mechanism which enhances the accuracy of generating corresponding 3D motions based on textual descriptions. AMD [158] achieves efficient conversion from text to 3D motion by fusing the original diffusion process with the reference diffusion process.
@@
In the Audio domain, Koizumi et al. [43] utilized an LLM, incorporating encoded dense features in the attention module to guide the generation of audio captions. Re-AudioLDM [159] utilizes distinct encoders to extract deep features from text and audio, which are then integrated into the attention mechanism of its Latent Diffusion Model (LDM).
@@
For video captioning, R-ConvED [48] uses a convolutional encoder-decoder network to process retrieved video-sentence pairs with an attention mechanism, generating hidden states to produce captions. CARE [160] introduces a concept detector to produce concept probabilities, and incorporates concept representations into a hybrid attention mechanism. EgoInstructor [49] employs gated-cross attention to integrate textual inputs with encoded video features, enhancing the relevance and coherence of the generated captions for egocentric video content.
Finally, Latent Representation-based RAG is adaptable to various modalities and tasks. It obtains the hidden states of retrieved data, enabling seamless integration between retrievers and generators. However, additional training is required to align the latent space. Within this paradigm, we have the flexibility to design intricate and novel algorithms that effectively combine the information from retrieved contents.
@@ III-A3 Logit-based RAG
In logit-based RAG, generative models integrate retrieval information through logits during the decoding process. Typically, the logits are combined through simple summation or models to compute the probabilities for step-wise generation.
@@
The kNN-LM [37] model integrates a pre-trained neural language model with the k-nearest neighbor search. It employs the pre-trained model to generate a list of candidate words and their probability distribution, while simultaneously performing retrieval from a data repository to find the k most relevant neighbors based on the current context, thus enhancing the output of the original language model. The innovation at the core of this model lies in its ability for dynamic retrieval of information from a broad text corpus, significantly improving the accuracy and relevance of its predictions, particularly in addressing rare patterns and adapting to various fields. He et al. [38] introduced a new framework that is predicated on performing retrieval operations only when necessary, aimed at enhancing the inference efficiency of the kNN-LM model through an adaptive retrieval. This framework accelerates the model’s inference speed by training a retrieval adapter, which automatically identifies and eliminates unnecessary retrieval actions in certain scenarios. This method allows the model to dynamically decide on the necessity of retrieval based on the current context, thereby balancing the trade-off between performance and efficiency, and substantially increasing inference speed while maintaining model performance.
@@
Unlike previous methods that only merge memories during the testing time, TRIME [161] achieves memory merging during both training and testing phases, treating in-batch examples as accessible memory. It leverages new data batching and memory construction techniques to effectively utilize external memory. It employs BM25 scores to pack paragraphs with high lexical overlap into the same batch, constructing the training memory to further optimize model performance. NPM [162] is a non-parametric masked language model comprised of an encoder and a reference corpus. Unlike traditional models that apply a softmax over a finite vocabulary, NPM models a non-parametric distribution over the corpus. The encoder’s role is to map phrases from the corpus into fixed-size vectors, filling in [MASK] by retrieving the phrase most similar to the masked position.
@@
Beyond Text, other modalities, such as code and Image, also leverage logit-based RAG. For code-to-text conversion task, Rencos [121] generates multiple summary candidates in parallel from the retrieved code. It then normalizes these candidates using edit distance and calculates the final probability to select the summary output that best matches the original code. In code summarization task, EDITSUM [138] enhances the quality of summary generation by integrating prototype summaries at the probability level. For text-to-code tasks, the kNN-TRANX [163] model employs a combination of a confidence network and meta-knowledge to merge retrieved code fragments. It utilizes a seq2tree structure to generate target code that closely matches the input query, thereby increasing the accuracy and relevance of code generation. In image captioning tasks, MA [164] combines an attention-based encoder-decoder, using the image encoder to extract visual features to construct the semantic part, and decodes it word by word with the information retrieved. MA interpolates between two distributions generated by the caption decoder and the memory-augmented module to determine the distribution of the next word.
In conclusion, Logit-based RAG effectively leverages historical states (or other data sources) to infer the current state and combines information at the logit level. This approach is well-suited for sequence generation tasks. It couples retrieval and generation, primarily requiring training of the generator. Novel approaches can be designed to effectively utilize the obtained probability distributions and adapt to subsequent tasks.
@@ III-A4 Speculative RAG
Speculative RAG seeks opportunities to use retrieval instead of pure generation, aiming to save resources and accelerate response speed. REST [32] replaces the small models in speculative decoding [165] with retrieval, enabling the generation of drafts. GPTCache [39] addresses the issue of high latency when using the LLM APIs by building a semantic cache for storing LLM responses. COG [166] decomposes the text generation process into a series of copy-and-paste operations, retrieving words or phrases from the documents instead of generation. Cao et al. [167] proposed a new paradigm to eliminate the dependence of the final result on the quality of the first-stage retrieved content, replacing generation with directly retrieved phrase level content.
In conclusion, Speculative RAG is currently primarily applicable to sequential data. It decouples the generator and the retriever, enabling the direct use of pre-trained models as components. Within this paradigm, we can explore a wider range of strategies to effectively utilize the retrieved content.
@@ III-B RAG Enhancements
In this section, we introduce methods which enhance the performance of a constructed RAG system. We categorize existing methods into 5 groups based on their enhancement targets: input, retriever, generator, result, and the entire pipeline.
@@ III-B1 Input Enhancement
The input, initially fed into the retriever, significantly impacts the final outcome of the retrieval stage. In this section, we introduce two methods for input enhancement: query transformation and data augmentation.
@@
Query transformation can enhance the result of retrieval by modifying the input query.
Query2doc [168] and HyDE [169] use the original query to generate a pseudo document, which is later used as the query for retrieval. The pseudo document contains richer relevant information, which helps to retrieve more accurate results. TOC [170] employs recursive retrieval augmented clarification on ambiguous questions to construct a tree of disambiguated questions, thereby generating comprehensive answers to these ambiguous questions. During the construction process of this tree structure, TOC utilizes a self-verification pruning method to ensure the factual relevance of each node.
@@
Data augmentation improves data before retrieval, including techniques such as removing irrelevant information, eliminating ambiguity, updating outdated documents, synthesize new data, etc.
Make-An-Audio [44] uses captioning and audio-text retrieval to generate captions for language-free audio to mitigate data sparsity, and adds random concept audio to improve the original audio. LESS [171] strategically selects an optimal dataset for downstream tasks by analyzing gradient information. It aims to maximize the dataset’s impact on fine-tuning the model’s performance in response to instructional prompts. ReACC [132] employs data augmentation (including renaming and dead code insertion) to pre-train the code retrieval model.
@@ III-B2 Retriever Enhancement
The retrieval process is crucial in RAG systems. Generally, the better the retrieved content quality, the easier it is to stimulate the ability of LLMs in-context learning as well as other generators and paradigms. The worse the content quality, the more likely it is to cause model hallucinations. Therefore, in this section, we will discuss how to efficiently improve the effectiveness of the retrieval process.
Recursive Retrieval
Recursive retrieval is to perform multiple searches to retrieve richer and higher-quality contents.
ReACT [172] uses Chain-of-Thought (CoT) [173] to break queries down for recursive retrieval and provide richer information. RATP [174] uses the Monte-Carlo Tree Search (MCTS) to perform multiple simulations, identifying optimal retrieval content. This content is then integrated into a template and sent to the generator for final production.
@@
Chunk optimization refers to adjusting chunk size for improved retrieval results. Sentence-window retrieval [175] is an efficient approach that enhances retrieval by fetching small chunks of text and returning a window of relevant sentences surrounding the retrieved segment. This method ensures that the context before and after the targeted sentence is included, providing a more comprehensive understanding of the retrieved information. Auto-merge retrieval is another advanced RAG method of LlamaIndex [175] which organizes the document in a tree-like structure, with the parent node containing the content of all children nodes. For example, articles and paragraphs, as well as paragraphs and sentences, all follow a parent-child relationship. In the retrieve process, fine-grained search for children nodes ultimately returns the parent node, effectively providing richer information. To address the lack of contextual information, RAPTOR [176] employs recursive embedding, clustering, and summarization of text chunks until further clustering becomes infeasible, thereby constructing a multi-level tree structure.
Retriever Finetuning
@@
As a core component in the RAG system, the retriever plays a crucial role in the entire system operation process. An effective embedding model can cluster semantically similar content in vector space, enhancing the retriever’s capability to provide valuable information for the subsequent generator, thus boosting the RAG system’s efficiency. Hence, the proficiency of the embedding model [177, 178, 179, 180] is vital for the RAG system’s overall effectiveness.
@@
In addition, for embedding models that already have good expression power, we can still finetune them using high-quality domain data or task related data to improve their performance in specific domains or tasks. REPLUG [127] treats LM as a black box and update the retriever model based on the final results. APICoder [129] finetunes the retriever with python files and api names, signature, description. EDITSUM [138] finetunes the retriever to decrease the jaccard distance between summaries after retrieval. SYNCHROMESH [122] adds tree distance os ASTs in the loss and uses Target Similarity Tuning to finetune the Retriever. R-ConvED [48] finetunes the Retriever with the same data as generator. Kulkarni et al. [181] applied infoNCE loss to finetune the Retriever.
Hybrid Retrieval
Hybrid retrieve denotes the concurrent employment of a diverse array of retrieval methodologies or the extraction of information from multiple distinct sources.
@@
RAP-Gen [182] and ReACC [132] use both dense retriever and sparse retriever to improve the quality of retrieval. Rencos [121] uses sparse retriever to retrieve similar code snippets on syntactic-level and usse dense retriever to retrieve similar code snippets on semantic-level. BASHEXPLAINER [139] first uses dense retriever to capture semantic information and then uses sparse retriever to acquire lexical information. RetDream [50] first retrieves with text and then retrieves with the image embedding. CRAG [183] has designed a retrieval evaluator to assess the relevance of retrieved documents to the input query, triggering three types of retrieval actions based on varying confidence levels: if deemed correct, it directly uses the retrieval results for Knowledge Refinement; if incorrect, it resorts to Web Search; and if ambiguous, it combines both approaches. To enhance performance in question-and-answer tasks, Huang et al. [184] introduced two metrics, DKS(Dense Knowledge Similarity) and RAC(Retriever as Answer Classifier), during the retrieval phase. These metrics account for both the pertinence of the answers and the applicability of the underlying knowledge. UniMS-RAG [185] introduces a novel kind of token, termed as the “acting token”, which determines the source from which to retrieve information.
@@
The Rerank technique refers to reordering the retrieved content in order to achieve greater diversity and better results. Re2G [186] applies a re-ranker [187] model after the traditional retriever. The effect of the re-ranker model is to re-rank retrieved documents, the purpose of which is to reduce the impact of information loss caused by compressing text into vectors on the quality of retrieval. AceCoder [188] reranks the retrieved programs with a selector to reduce redundant programs and obtain diverse retrieved programs. XRICL [189] uses a distillation-based exemplar reranker after retrieval. Rangan [190] employs the Quantized Influence Measure, assessing statistical biases between a query and a reference to evaluate the similarity of data subsets and rerank retrieval results. UDAPDR [191] uses LLMs to cost-effectively generate synthetic queries that train domain-specific rerankers, which then apply multi-teacher knowledge distillation to develop a cohesive retriever. LLM-R [192] iteratively trains the retriever, using feedback from a frozen LLM to rank candidate documents and train the reward model. The retriever is further trained based on knowledge distillation. Each iteration of training builds upon the retriever trained in the previous cycle, facilitating iterative optimization in subsequent rounds.
@@
Retrieval Transformation involves rephrasing retrieved content to better activate the generator’s potential, resulting in improved output.
FILCO [193] effectively filters out irrelevant content from the retrieved text chunk, leaving only the precise supporting content. This process simplifies the task for the generator, making it easier to predict the correct answer. FiD-Light [194] initially employs an encoder to convert the retrieved content into a vector, which it then compresses, resulting in a substantial reduction of latency time. RRR [195] integrates the current query with the top-k document in each round through a template, and subsequently restructures it via a pre-trained LLMs(GPT-3.5-Turbo etc.).
@@
In addition to the above optimization methods, there are also some other optimization methods for the retrieve process. For example, Meta-data filtering  [196] is a method to help processing retrieved documents which uses metadata (such as time, purpose, etc.) to filter the retrieved documents for better results. GENREAD [197] and GRG [198] introduce a novel approach where the retrieval process is supplanted or improved by prompting a LLM to generate documents in response to a given question.